OpenStudy (anonymous):

z^2+iz-6z+10 What are the solutions for z? Note i is the imaginary number...

OpenStudy (anonymous):

This one has two complex roots; z = 2 + i, and z = 4 - 2i :)

OpenStudy (anonymous):

Thank you! Could i ask how you solved it?